Subject: Minutes for January 6, 1997 User's Committee Meeting

From: strauss@astro.Princeton.EDU

Submitted: Wed, 8 Jan 1997 14:07:44 -0500 (EST)

Message number: 113 (previous: 112, next: 114 up: Index)

   Minutes of APO 3.5m User's Committee Phone Conference
		  Monday, January 6, 1997

   Attending: Bruce Gillespie, Ed Turner (Chair), Michael Strauss
(taking minutes), Rene Walterbos, Julie Lutz

  Given the small number of people attending this meeting, we decided
to have an abbreviated meeting, and spent most of the time discussing
the telescope shutdown, which is a few days from ending.  This
shutdown was to realuminize all the mirrors, and to replace the
enclosure wheels. 

Gillespie: The APO staff is putting everything back together today,
rough alignment in progress, with full alignment and pointing model
(probably requiring 1.5 nights of clear skies) starting tomorrow
night.  The earliest the telescope could be given back to the
scientists for observations is thus Wednesday or Thursday night, but the
mountaintop is in the middle of a blizzard right now, and so there may
be further delays. 

  One enclosure wheel was replaced about 9 months ago; during the
shutdown, the other three enclosure wheels were also replaced.  It was
anticipated that this would be more difficult, given the fact that
some of these wheels had drive motors on them, but all went very
smoothly.  Many kudos to Jon Davis and his group, who were able to do
the last three wheels without help from L&F. 

  All three mirrors have been realuminized, at KPNO (primary) and Sac
Peak (secondary, tertiary); the reflectivity from each as a function
of wavelength looks beautiful.  Kudos to Mark Klaene to making this
all work. 

  When the primary was taken out, it was noticed that some of the
ventilation tubes that force air through the cells in the back of the
primary were loose; the epoxy that held them in place had come loose.
It turns out that fixing them properly would have been a major job,
and therefore the decision was made simply to take the loose ones out
altogether pending a redesign.  If that turns out to degrade the
seeing, then something will have to be jerry-rigged to put them back
in. It is expected that a better system for attaching the tubes to the
primary mirror cell will be developed and implemented at some future

  When the secondary was taken out, it was noticed that the flex-joints for
the whiffletrees connected to the secondary servos were all broken.
These have now been replaced.  The secondary support
system may need to be replaced with a different design in the future.

  It was also noticed that the piezoelectric stacks for the secondary,
which are in place for future tip-tilt work, had all failed, and will
need to be repaired, and have been replaced by dummy cylinders for the
time being. 

  As soon as there is some clear weather, new throughput measurements will
be taken.  Everyone is hopeful that with the realuminization, the
throughput will have gone up substantially. 

  The primary mirror support system is in the process of being
redesigned for higher bandwidth and better stability, with the aim to
reduce low frequency mirror oscillations.  The new system should be in
place on a 1-2 month timescale, and will not require any major
telescope shutdown while it is in progress.

  Thermal sensors in the mirror have been replaced by Alan Diercks. 

  During the shutdown, Jon Brinkman and Alan Uomoto worked on the DIS
electronics, and are close to eliminating a source of AC pickup in the
CCD's; they are waiting on one new part. 

  Charles Corson, previously of KPNO, is joining the APO 3.5m staff. 

  Progress has been made on finding a vendor for polishing the new
secondary; we are talking to several different vendors.  We should be
closer to making a decision in about 3 weeks.

  There will be a four-night block of time given over to engineering
in late January.  It has been scheduled for installation of the new
guider and to work on the primary support system.  We will make a
larger effort in the future to send out information to the user
community about what actually happens during these engineering runs. 

  Last meeting's minutes are approved. 

  Next meeting, February 10, 12:30 PM. 

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 113 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to